# Effect of ultralight dark matter on compact binary mergers

### v1

@@ I Introduction
However, the significance of studying compact object environments extends beyond these aspects. It can also reveal crucial information about a variety of astrophysical phenomena, including the presence of a third body—such as exoplanets or stars—orbiting these compact objects [11, 12]. Additionally, such investigations can shed light on the existence and distribution of dark matter (DM) [13, 14, 15]. In this work, we focus specifically on exploring the presence and properties of DM based on current observational data.
@@
The fact that our universe entails more invisible matter than visible ones had its origin in the observations involving rotation curves of galaxies, where it was observed that the almost-constant rotation curves can be beautifully explained by postulating the existence of such invisible or, dark matter [22]. Besides, several other clinching astrophysical evidences, e.g., bullet cluster, dwarf spheroidal galaxies, gravitational lensing have been put forward in support of the DM hypothesis [23, 24, 25]. Similarly, in the context of cosmology as well, the matter content of our universe requires the existence of DM, whose abundance is at least five times larger than those in the standard model of particle physics [26]. In particular, according to the current paradigm of cosmology, DM plays a crucial role in the formation of structures in our universe, as they are the first one to collapse into structures called DM halos, into which visible matter falls, leading to individual galaxies. Observational constraints from structure formation [27], Lyman-$\alpha$ measurements [28] and lensing [29] restrict the composition of DM at length scales $\gtrsim 100$ Mpc to be cold (CDM). In other words, DM behaves as a pressure-less fluid at these scales. This gives rise to the well-known $\Lambda$CDM paradigm, the standard model of cosmology. However, at small scales, the CDM hypothesis faces serious challenges, the ‘Core-Cusp problem’ [30], the ‘Galactic Morphology problem’ [31], and the ‘Missing satellites problem’ [32] are to name a few of them. Side by side, the observational constraints on the nature of DM are also much weaker at galactic length scales of $\sim 10-100$ Kpc.
One way to circumvent this difficulty, is to assume that DM is no longer cold in these regions, rather they form a condensate. This is where, the Ultra-Light DM (ULDM) plays a pivotal role [33] and can alleviate some of the challenges faced by the CDM scenario. In this formulation, DM is thought to be composed of light bosons, which forms a condensate at galactic length scales, thereby deviating from the CDM hypothesis, while smoothly joins with CDM at larger length scales, without violating any observational constraints [33]. Moreover, such light bosons have also been shown to arise from simple extensions of the standard model for particle physics, e.g., models involving axions [34]. All of these make ULDM models a viable alternative to the CDM at the galactic scale.
@@
[…]
@@ II.1 Framework
[…]
@@
Since we observe the GWs emitted from such systems at infinity, our frame of reference coincides with the rest frame of the scalar field. Hence all the expressions ought to be written in the rest frame of the scalar field. In this frame, the objects move in their binary orbits through a homogenous ULDM density. This is simply achieved by the following transformations: $\dot{M}^{\prime}=(\dot{M}/\gamma)$ and $F^{\prime}_{i}=F_{i}+\dot{M}v_{i}$. Then, it turns out (see [39] for the detailed calculation) that the accretion rate of an object of mass $M$ and angular momentum $J$ moving through the ULDM medium with an orbital speed $v$, in the rest frame of the DM, is given by,
@@ II.2 Implementation in binary inspiral
$-\frac{1}{2}\frac{d}{dt}\left(\frac{m_{\rm r}M_{\rm T}}{a}\right)=-\frac{32}{5}\frac{m_{\rm r}^{2}M_{\rm T}^{3}}{a^{5}}-\frac{4\pi\hbar n}{\mu}\sum_{i=1}^{2}\eta^{2}_{i}v_{i}\Delta\Psi_{i}\,,$ (10)
@@ III.1 Individual Binaries
We show the effect of the presence of ULDM on the merger timescale of individual binaries in this section. For that we start with a chosen mass configuration of binary main sequence stars as progenitors and track their merger status by evolving Eq. (II.2) to obtain the binary separation $a$ with time. To do this consistently, we require a proper transformation function when a progenitor star reaches its stellar endpoint. We implement this by a simple mass-transfer function [42] to determine the mass of the evolutionary remnant of the main sequence star. In reality, such endpoints of massive main-sequence stars are usually accompanied by novas, which are highly model dependent and is likely to add a further layer of uncertainty to our baseline model, but will not significantly contribute to an understanding of the effects of ULDM on binary merger. Nevertheless, we need to implement a model to facilitate mergers and not disrupt them. To do this, we adopt a method similar to the evolution method previously implemented in [43] with the notable exception that our method does not model for any additional kick at the endpoint of main-sequence evolution. Thus we will potentially be biased towards higher merger rate. To determine the shift in the orbital separation $a$ due to the stellar endpoint, we consider the variation in the angular momentum $L\coloneqq m_{\rm r}a^{2}\Omega$ and use the Kepler’s third law $\Omega^{2}a^{3}=M_{\rm T}$, to obtain,
$\left(\frac{\delta a}{a}\right)$ $=2\left(\frac{\delta L}{L}\right)-2\left(\frac{\delta m_{\rm r}}{m_{\rm r}}\right)-\left(\frac{\delta M_{\rm T}}{M_{\rm T}}\right)\,,$ (11)
$\left(\frac{\delta\Omega}{\Omega}\right)$ $=-3\left(\frac{\delta L}{L}\right)+3\left(\frac{\delta m_{\rm r}}{m_{\rm r}}\right)+2\left(\frac{\delta M_{\rm T}}{M_{\rm T}}\right)\,,$ (12)
@@
where $\delta x$ represents the change in a quantity $x$ before and after the stellar endpoint. We can then eliminate $(\delta L/L)$ from Eq. (11) to obtain $3(\delta a/a)=(\delta M_{\rm T}/M_{\rm T})-2(\delta\Omega/\Omega)$. From the mass-transfer function implementation of [42] the final mass is less than the initial mass because a nova usually involves ejection of mass. Therefore, it follows that $(\delta M_{\rm T},\delta m_{\rm r})<0$. We further assume that $(\delta\Omega/\Omega)\geq 0$, which ensures that the binary separation does not increase across the stellar endpoint. Further, from the above expressions it follows that $|\delta a/a|\geq(1/3)|\delta M_{\rm T}/M_{\rm T}|$. From this, we finally make our choice of $\delta a$ across stellar endpoint as
@@
III.2 Evolution of a binary population of progenitors
@@
Consequently, we will assume a uniform IMF in order to be maximally agnostic while continuing to assume no effects of eccentricity or BH spin. Moreover, our evolution also does not factor other astrophysical effects like Roche overflows, stellar winds or common envelope binary evolution. We remark that factoring in all of these components that we have left out require a fully numerical population synthesis, which is beyond the scope of this project. The choice of the uniform IMF means that we also need to choose appropriate endpoints in the progenitor stellar-mass distribution. We choose a lower cut-off of the progenitor mass, such that the IMF will have masses satisfying, $M_{i}^{\rm P}\geq 10M_{\odot}$ based on physical considerations. This implies that the stars having masses below this cut-off will not produce compact remnants. Masses of stars are also bounded from above to $M_{i}^{\rm P}\leq 100M_{\odot}$ from considerations of hydrostatic equilibrium. For our purposes in this work, we choose a slightly lower value for the upper limit of the progenitor masses, namely, $M_{i}^{\rm P}\leq 80M_{\odot}$. We argue that the bias produced through this choice will be negligible (compared to other approximations) given that IMF generically exhibits rapid decay for stellar masses above a few solar mass.
@@
We now return to the question of the final state of main sequence star progenitors, all of which evolve in the presence of ULDM, while starting at a fixed redshift. We present the answer to this question in Fig. 3, where we show that the end states of the evolution for progenitor binaries starting from $z=0.5$. Here and henceforth, we adopt the convention that for the progenitors $M_{2}^{\rm P}\leq M_{1}^{\rm P}$, such that the mass-ratio $q^{\rm P}\leq 1$. Following which, in Fig. 3, we show four different panels correspond to four different ambient densities of the ULDM environment. For each panel, we show the mass ratio of the progenitors $q^{\rm P}$ on the x-axis and the total mass of the binary configuration $M_{\rm T}^{\rm P}=M_{1}^{\rm P}+M_{2}^{\rm P}$ on the y-axis. Following our reasoning above, the grey shaded regions represent configurations that are off-limits to our interest, i.e., with progenitors masses being too high, or, too low. It is expected that the merger timescales for the binary systems will not all be the same, and thus we would get a combination of merged and unmerged binary populations, depending on the DM density in the environment. In Fig. 3, the red line acts as a separatrix, which separates the configuration space into merged and unmerged endpoints for the binaries. The redshifts for the merged population are shown by the green-yellow shade, while for the unmerged populations we depict the binary separation at $z=0$ by the blue shade.
@@
Table 1: Table showing the fraction of progenitors that have merged at $z=0$, starting their orbital evolution at $z=0.5$, with an increase in the ULDM density $\rho$. Note that the conservation between the unit for the DM density used in this work with the density in $\textrm{gm}/\textrm{cm}^{3}$ is: $1~{}\text{GeV}/\textrm{cm}^{3}=1.78\times 10^{-24}~{}\text{gm}/\textrm{cm}^{3}$.
@@ IV Effects on Astrophysical Merger Rates
The results of the previous section have firmly established that presence of an ULDM environment should have an effect on the population statistics of mergers. However, we only considered progenitor sources at a single redshift $z=0.5$. A complete picture can be obtained if we can repeat the above exercise, but factoring in sources from a given range of redshifts. Additionally, we recall that all redshifts will not contribute equally to the generation of compact binary progenitors. This is because the comoving density of main-sequence stars is not constant with the redshift, and has a functional dependence which is well-known in the literature as the Star Formation Rate (SFR). In our analysis in this section, we have used the SFR function, as proposed in Madau & Dickinson [52]. At this outset, we would like to emphasize that the SFR simply provides the total mass of the stars produced at a given redshift per comoving volume. One needs to convolve it with the IMF in order to get the number density of the stars. However in our case this will just be a constant factor, given that our IMF is assumed to be uniform. Subsequent multiplication with the comoving volume results into a function for the number of sources in the redshift bin $z$ and $z+dz$, which can be observed per unit observer’s time. This function, representing number of binary progenitors per unit redshift per unit time, is given by
@@
where $\mathcal{R_{\psi}}(z)$ is the SFR at redshift $z$, and $(dV_{\rm c}/dz)$ is arising from the cosmological comoving volume. The factor $(1+z)^{-1}$ appears because of the difference in the time intervals at the source and at the observer. Given Eq. (15), we are now in a position to compute the rate of binary compact mergers for our underlying astrophysical model. To compute the ‘merger-rate density’ we need to first evolve all sources at a given $z$, given by Eq. (15), either to their merger or to $z=0$. We then bin the population of all the merged events according to their merger redshifts, which gives for us the number $(d^{2}N_{\rm m}/dz_{\rm m}dt)$, depicting the merger events between redshifts $z_{\rm m}$ and $z_{\rm m}+dz_{\rm m}$, observed per unit observer’s time. Here, $z_{m}$ represents the merger redshifts. Dividing the above by the comoving volume factor at $z_{m}$, we finally arrive at our estimate for the merger rate density,
$\mathcal{R}_{\rm m}(z_{\rm m})\sim\frac{d^{2}N_{\rm m}}{dz_{\rm m}dt}\left(\frac{dz_{\rm m}}{dV_{\rm c}}\right)~{}.$ (16)
We can now compare our results and match our estimates of the merger rate $\mathcal{R}_{\rm m}(z_{\rm m})$ to the inferred rate $\mathcal{R}_{\rm m}^{\mathrm{inf}}(z_{\rm m})$ from the LIGO-Virgo observations. But before we do that a word of caution is in order. We reiterate that our baseline model does not contain information about the source IMF and its dependence on $z$. It also does not consider inputs from many dissipative astrophysical effects (like Roche overflows, common envelope evolutions etc.). Therefore, it is natural to expect overestimation for our calculated merger rates. We should also remark that at current sensitivity, GW observations can only observe the ‘local’ merger rate density $\mathcal{R}_{\rm m}^{\mathrm{obs}}(0)$, i.e., the merger rate density at $z=0$. The overestimation can nevertheless be removed by an overall rescaling of the number of sources such that the rescaled estimates of the rate $\mathcal{R}_{\rm m}(z_{\rm m})$ obey observational bounds at $z_{\rm m}=0$. In other words, such a rescaling will make $\mathcal{R}_{\rm m}(0)$ lie within the error bars of $\mathcal{R}_{\rm m}^{\mathrm{obs}}(0)\equiv R_{0}$. Subsequently, following this standard procedure in the population synthesis literature (see, for example [53]) we make use of appropriate rescaling functions to bring the value of $\mathcal{R}_{\rm m}(0)$ within the GWTC-3 bounds [54].

### v2

@@ I Introduction
However, the significance of studying compact object environments extends beyond these aspects. It can also reveal crucial information about a variety of astrophysical phenomena, including the presence of a third body—such as exoplanets or stars—orbiting these compact objects [57, 51]. Additionally, such investigations can shed light on the existence and distribution of dark matter (DM) [46, 26, 14, 19, 22, 15]. In this work, we focus specifically on exploring the presence and properties of DM based on current observational data.
@@
[…]
One way to circumvent this difficulty, is to assume that DM form a condensate. This is where, the models of Ultra-Light DM (ULDM) play a pivotal role [29] and can alleviate some of the challenges faced by the CDM scenario. In this formulation, DM is thought to be composed of light bosons, which forms a condensate at sub-halo length scales, while smoothly joins with CDM at larger distances, without violating any observational constraints [29]. Moreover, such light bosons have also been shown to arise from simple extensions of the standard model for particle physics, e.g., models involving axions [31]. All of these make ULDM models a viable extension of the CDM at the galactic scale.
@@
Conventions — Throughout this paper we will use geometric units, namely $G=c=1$. Note that we do not use $\hbar=1$. Also, we work in four spacetime dimensions, with mostly positive signature convention, such that the flat metric becomes $\eta_{\mu\nu}=\textrm{diag.}(-1,+1,+1,+1)$.
@@ II.1 Framework
$\dot{M}$ $=\int_{\infty}r^{2}T_{tr}d\Omega_{2}=\rho A_{\rm h}\left(\frac{e^{\pi\eta}\pi\eta}{\sinh(\pi\eta)}\right)\left(\frac{1-R}{1+R}\right)~,$ (3)
$F^{i}$ $=\int_{\infty}r^{2}T^{i}_{r}d\Omega_{2}=-\frac{4\pi\rho M^{2}}{v^{2}}(1+v^{2})^{2}\mathcal{D}\zeta^{i}-\dot{M}v^{i}~.$ (4)
@@
The above analysis was performed in the frame of the compact objects, while the observation is being made at asymptotic infinity. Since the DM is static in the observer’s frame of reference, we need to make a transformation from the frame of the BH to the frame of the observer. This is simply achieved by the following transformations: $\dot{M}^{\prime}=(\dot{M}/\gamma)$ and $F^{\prime}_{i}=F_{i}+\dot{M}v_{i}$. Then, it turns out (see [49] for the detailed calculation) that the accretion rate of an object of mass $M$ and angular momentum $J$ moving through the ULDM medium with an orbital speed $v$, in the rest frame of the DM, is given by,
@@ II.2 Implementation in binary inspiral
$\frac{1}{2}\frac{d}{dt}\left(\frac{m_{\rm r}M_{\rm T}}{a}\right)=-\frac{32}{5}\frac{m_{\rm r}^{2}M_{\rm T}^{3}}{a^{5}}-\frac{4\pi\hbar n}{\mu}\sum_{i=1}^{2}\eta^{2}_{i}v_{i}\Delta\Psi_{i}\,$, (10)
@@ III.1 Individual Binaries
We show the effect of the presence of ULDM on the merger timescale of individual binaries in this section. For that we start with a chosen mass configuration of binary main sequence stars as progenitors and track their merger status by evolving Eq. (II.2) to obtain the binary separation $a$ with time. To do this consistently, we require a proper transformation function when a progenitor star reaches its stellar endpoint. We implement this by a simple mass-transfer function [32] to determine the mass of the evolutionary remnant of the main sequence star. In reality, such endpoints of massive main-sequence stars are usually accompanied by novas, which are highly model dependent and is likely to add a further layer of uncertainty to our baseline model, but will not significantly contribute to an understanding of the effects of ULDM on binary merger. Nevertheless, we need to implement a model to facilitate mergers and not disrupt them. To do this, we adopt a method similar to the evolution method previously implemented in [23] with the notable exception that our method does not implement a constant ejecta velocity to model the kick imparted to the stellar remnants. Rather, we let the ejecta velocity to depend on the initial masses of the stellar progenitors and also on the associated mass loss. In this case, the kick imparted to the remnant is likely to cause a sudden change to the angular momentum, depending on various parameters of the system. The change in angular momentum also causes a change in the angular velocity of the remnants. These changes can be determined through the following equations obtained from $L\coloneqq m_{\rm r}a^{2}\Omega$ and Kepler’s third law $\Omega^{2}a^{3}=M_{\rm T}$.
$\left(\frac{\delta\Omega}{\Omega}\right)$ $=-3\left(\frac{\delta L}{L}\right)+3\left(\frac{\delta m_{\rm r}}{m_{\rm r}}\right)+2\left(\frac{\delta M_{\rm T}}{M_{\rm T}}\right)\,.$ (11)
@@
Here $\delta x$ represents the change in a quantity $x$ before and after the stellar endpoint. From the mass-transfer function implementation of [32] the final mass is less than the initial mass because a nova usually involves ejection of mass. Therefore, it follows that $(\delta M_{\rm T},\delta m_{\rm r})<0$. Using these results in Eq. (11), we notice that only if $\delta L<0$, one might have $\delta\Omega>0$. Then to connect with the binary separation at the stellar endpoint, we use Kepler’s third law again to obtain,
$3(\delta a/a)=(\delta M_{\rm T}/M_{\rm T})-2(\delta\Omega/\Omega)$ (12)
This suggests that whenever $(\delta\Omega/\Omega)\geq 0$, the binary separation always decreases at stellar endpoint. Therefore we work with precisely those events where $(\delta\Omega/\Omega)\geq 0$ is obeyed at mass transfer, which in turn ensure that the binary separation never increase across the stellar endpoint. We note that the condition $\delta L<0$ implies that the angular momentum induced by the ejecta is opposite to that of the orbital angular momentum. On the other hand, the condition $\delta L\geq 0$ increases the binary separation and can lead to potential disruption of the system [11]. Hence on average, the number of binaries whose separation between the individual components decrease will be reduced by some factor. Luckily, this decrease in the number of such events does not affect our results because we rescale the merger events so that the local merger rate density matches observations (see Sec  IV).
Further, from the above expressions it follows that $|\delta a/a|\geq(1/3)|\delta M_{\rm T}/M_{\rm T}|$. From this, we finally make our choice of $\delta a$ across stellar endpoint as
@@
III.2 Evolution of a binary population of progenitors from a given redshift
@@
Consequently, we will work with both uniform as well as non-unform (Kroupa) IMFs [43] in our analysis. The uniform IMF, though probably unphysical still allows us to be maximally agnostic, while the Kroupa model is motivated by the stellar observations from the galactic plane. Also, we will continue to assume no effects of eccentricity in the binary orbit or BH spin in our analysis. Moreover, our evolution does not include other astrophysical effects, like Roche overflows, stellar winds or common envelope binary evolution. We remark that factoring in all of these components that we have left out require a fully numerical population synthesis, which is beyond the scope of this work. Therefore, what we are assuming is that the dynamical friction and energy loss by GWs are the only dissipative components in the dynamics. So, any constraint on the density of the ULDM from our analysis should be treated only as an upper limit.
Let us now focus on our choice of lower and upper cut-off for the IMFs. It is reasonable to expect that progenitor stars which have masses below a limit cannot produce compact remnants. Keeping this in mind we choose a lower cut-off of the progenitor masses given by $M_{i}^{\rm P}\geq 10M_{\odot}$. Then the use of the uniform IMF requires us to also choose an appropriate upper bound in the progenitor stellar-mass distribution. Masses of stars are bounded from above to $M_{i}^{\rm P}\leq 100M_{\odot}$ from considerations of hydrostatic equilibrium. For our purposes in this work, we choose a slightly lower value for the upper limit of the progenitor masses, namely, $M_{i}^{\rm P}\leq 80M_{\odot}$. A limitation of this mass cut-off could be that it might produce bias for the choice of uniform IMF. However, the bias for the case of the more realistic non-uniform IMF will be negligible, as that IMF function decays to zero at those high mass values.
@@
A final uncertainty which plays a role in determining the merger timescales is the choice of initial separation of the progenitor stars. However, there are still some physically motivated factors which help to narrow down the range of the initial separations. We note that the initial separation is bound from below if we demand no Roche overflow in the initial progenitor configuration. Consequently we assume the initial configuration of the binaries to always obey the no Roche overflow condition. We have implemented this by making the sum of the initial Roche lobes of the component stars $2.5$ times the sum of their radii. The rescaling of merger rate (see Sec  IV) ensures that the rate of mergers in our model is commensurate with observations even with our specific choice of initial separations. We now return to the question of the final state of main sequence star progenitors, all of which evolve in the presence of ULDM, while starting at a fixed redshift. We present the answer to this question in Fig. 3, where we show that the end states of the evolution for progenitor binaries starting from $z=0.5$. Here and henceforth, we adopt the convention that for the progenitors $M_{2}^{\rm P}\leq M_{1}^{\rm P}$, such that the mass-ratio $q^{\rm P}\leq 1$. Following which, in Fig. 3, we show four different panels correspond to four different ambient densities of the ULDM environment. For each panel, we show the mass ratio of the progenitors $q^{\rm P}$ on the x-axis and the total mass of the binary configuration $M_{\rm T}^{\rm P}=M_{1}^{\rm P}+M_{2}^{\rm P}$ on the y-axis. Following our reasoning above, the grey shaded regions represent configurations that are off-limits to our interest, i.e., with progenitors masses being too high, or, too low. It is expected that the merger timescales for the binary systems will not all be the same, and thus we would get a combination of merged and unmerged binary populations, depending on the DM density in the environment. In Fig. 3, the red line acts as a separatrix, which separates the configuration space into merged and unmerged endpoints for the binaries. The redshifts for the merged population are shown by the green-yellow shade, while for the unmerged populations we depict the binary separation at $z=0$ by the blue shade.
@@
Table 1: Table showing the fraction of progenitors that have merged at $z=0$, starting their orbital evolution at $z=0.5$, with an increase in the ULDM density $\rho$. Note that we use a value of $m_{\Phi}=10^{-15}\text{eV}$ to be consistent with the ultralight condition $M\mu\ll 1$. The conservation between the unit for the DM density used in this work with the density in $\textrm{gm}/\textrm{cm}^{3}$ is: $1~\text{GeV}/\textrm{cm}^{3}=1.78\times 10^{-24}~\text{gm}/\textrm{cm}^{3}$.
@@ IV Effects on Astrophysical Merger Rates
The results of the previous section have firmly established that presence of an ULDM environment should have an effect on the population statistics of mergers. However, we only considered progenitor sources at a single redshift $z=0.5$. A complete picture can be obtained if we can repeat the above exercise, but factoring in sources from a given range of redshifts. Additionally, we recall that all redshifts will not contribute equally to the generation of compact binary progenitors. This is because the comoving density of main-sequence stars is not constant with the redshift, and has a functional dependence which is well-known in the literature as the Star Formation Rate (SFR). In our analysis in this section, we have used the SFR function, as proposed in Madau & Dickinson [45]. At this outset, we would like to emphasize that the SFR simply provides the total mass of the stars produced at a given redshift per comoving volume. One needs to convolve it with the IMF in order to get the number density of the stars. For the cases we have considered, the convolution integrals will be non-trivial for the Kroupa IMF and just a constant factor in the case of flat IMF. Subsequent multiplication with the comoving volume results into a function for the number of sources in the redshift bin $z$ and $z+dz$, which can be observed per unit observer’s time. This function, representing number of binary progenitors per unit redshift per unit time, is given by
@@
where $\mathcal{R_{\psi}}(z)$ is the SFR at redshift $z$, and $(dV_{\rm c}/dz)$ is arising from the cosmological comoving volume. The factor $(1+z)^{-1}$ appears because of the difference in the time intervals at the source and at the observer. Given Eq. (15), we are now in a position to compute the rate of binary compact mergers for our underlying astrophysical model. To compute the ‘merger-rate density’ we need to first evolve all sources at a given $z$, given by Eq. (15), either to their merger or to $z=0$. We then bin the population of all the merged events according to their merger redshifts, which gives for us the number $(d^{2}N_{\rm m}/dz_{\rm m}dt)$, depicting the merger events between redshifts $z_{\rm m}$ and $z_{\rm m}+dz_{\rm m}$, observed per unit observer’s time. Here, $z_{m}$ represents the merger redshifts. Dividing the above by the comoving volume factor at $z_{m}$ and multiplying by a factor of $1+z_{\rm m}$ to account for the redshift factor, we finally arrive at our estimate for the merger rate density,
$\mathcal{R}_{\rm m}(z_{\rm m})\sim\frac{d^{2}N_{\rm m}}{dz_{\rm m}dt}\left(\frac{dz_{\rm m}}{dV_{\rm c}}\right)\left(1+z_{\rm m}\right)~.$ (16)
We can now compare our results and match our estimates of the merger rate $\mathcal{R}_{\rm m}(z_{\rm m})$ to the inferred rate $\mathcal{R}_{\rm m}^{\mathrm{inf}}(z_{\rm m})$ from the LIGO-Virgo observations. But before we do that a word of caution is in order. We reiterate that our baseline model does not consider inputs from many dissipative astrophysical effects (like Roche overflows, common envelope evolutions etc.). It also does not contain information about other non-dissipative astrophysics like for example the $z$ dependence of the stellar IMF or the fraction of progenitors that get trapped in binary systems. Therefore, it is natural to expect an incorrect estimation for our calculated merger rates. We should also remark that at current sensitivity, GW observations can only observe the ‘local’ merger rate density $\mathcal{R}_{\rm m}^{\mathrm{obs}}(0)$, i.e., the merger rate density at $z=0$. The inaccuracy in the estimation of the merger rate can nevertheless be removed by an overall rescaling of the number of sources such that the rescaled estimates of the rate $\mathcal{R}_{\rm m}(z_{\rm m})$ obey observational bounds at $z_{\rm m}=0$. In other words, such a rescaling will make $\mathcal{R}_{\rm m}(0)$ lie within the error bars of $\mathcal{R}_{\rm m}^{\mathrm{obs}}(0)\equiv R_{0}$. Subsequently, following this standard procedure in the population synthesis literature (see, for example [17]) we make use of appropriate rescaling functions to bring the value of $\mathcal{R}_{\rm m}(0)$ within the GWTC-3 bounds [5].